OpenStudy (anonymous):

Samir drove 50 km on 6 L of gasoline. Which of the following would be an equation to help you find out how far he can travel on a full tank of 40 L? choice are in comments below

OpenStudy (anonymous):

\[ \frac{ 50 }{ 6 } =\frac{ 40 }{ x } \[\frac{ 50 }{ 6 } = \frac{ x }{ 40 }\] \[\frac{ 50 }{ 40 }= \frac{ x }{ 6 }\] \[\frac{ 40 }{ 50 }=\frac{ 6 }{ x }\]

OpenStudy (anonymous):

@Nurali could you help me?

OpenStudy (johnweldon1993):

You pretty much had it right...50 km on 6 L of gas so 40L would be x amount of km \[\large \frac{50}{6} = \frac{x}{40}\] You would multiply both sides of the equation by 40 to isolate and solve for 'x'

OpenStudy (anonymous):

ok thank you for the help!